please help me on evaluating this(it is log question)
16.4*9814

Answer by stanbon(75887)

16.4*9814
------
log[16.4*9814] = log(16.4)+log(9814)
= 1.2148 + 3.9918
---
= 5.2067
------
So, 16.4*9814 = 10^ 5.2067 is approximately 160,953.34
=============
Cheers,
Stan H.